# EqvAfford: SE(3) Equivariance for Point-Level Affordance Learning

### v1

@@ 2.1 Learning Affordance for Robotic Manipulation
For manipulation, it’s crucial to find suitable interaction points. Gibson proposed affordance – interaction opportunity. They leveraged observations as demonstrations, like videos showing agents interacting with scenarios, and trained models to learn what and how to interact. This kind of method has been used to learn possible contact locations[1] or grasp patterns[5]. However, using passive observations to learn possible interactions suffers from distribution shifts. [13] proposed an approach learning from interactions to obtain more informative samples for articulated object manipulation[3, 7, 6]. Recent works expanded affordance to more diverse scenarios and tasks, such as door [10] and articulated object [19, 11] manipulation, deformable object manipulation [20, 12, 22], clutter manipulation [9], language-guided manipulation [24], bimanual manipulation [27] and active exploration [14, 17].
@@ 2.2 Leveraging SE(3) Equivariant Representations for Robot Manipulation
Many works focused on building SE(3) equivariant representations from 3D inputs[2, 4, 8]. to enable robust generalizations to out-of-distribution inputs by ensuring inherent equivariance. This allows the model to automatically generalize to inputs at novel 6D poses, without implementing strong data augmentation. Prior works have applied equivariant representations on diverse robot manipulation tasks [25, 15, 16, 21, 26]. This work studies leveraging SE(3) equivariance for point-level affordance learning and thus complete articulated object manipulation.

### v2

@@ 2.1 Learning Affordance for Robotic Manipulation
For manipulation, it’s crucial to find suitable interaction points. Gibson proposed affordance – interaction opportunity. They leveraged observations as demonstrations, like videos showing agents interacting with scenarios, and trained models to learn what and how to interact. This kind of method has been used to learn possible contact locations[1] or grasp patterns[6]. However, using passive observations to learn possible interactions suffers from distribution shifts. [17] proposed an approach learning from interactions to obtain more informative samples for articulated object manipulation[10, 8]. Recent works expanded affordance to more diverse scenarios and tasks, such as articulated object manipulation [35, 25, 5], deformable object manipulation [16, 30, 28], clutter manipulation [14], bimanual manipulation [36], and zero-shot generalization to novel shapes [11, 12].
@@ 2.2 Leveraging SE(3) Equivariant Representations for Robot Manipulation
Many works focused on building SE(3) equivariant representations from 3D inputs[2, 3, 13, 18, 26, 21, 34, 15, 20]. to enable robust generalizations to out-of-distribution inputs by ensuring inherent equivariance. This allows the model to automatically generalize to inputs at novel 6D poses, without implementing strong data augmentation. Prior works have applied equivariant representations on diverse robot manipulation tasks [32, 22, 23, 29, 33, 24, 7, 9, 4, 19]. This work studies leveraging SE(3) equivariance for point-level affordance learning and thus complete articulated object manipulation.